OpenStudy (cookiemonster2412):

HELP!!! 12r^2=7r^2+20

OpenStudy (raffle_snaffle):

solve for r?

OpenStudy (cookiemonster2412):

yes

OpenStudy (shamil98):

12r^2 = 7r^2 + 20 5r^2 = 20 r^2 = 4 \[\ r = \pm 2\]